# CNN on ‘Top’: In Search of Scalable & Lightweight Image-based Jet Taggers

### v1

@@ Abstract
While Transformer-based and standard Graph Neural Networks (GNNs) have proven to be the best performers in classifying different types of jets, they require substantial computational power. We explore the scope of using a lightweight and scalable version of EfficientNet architecture, along with global features of the jet. The end product is computationally inexpensive but is capable of competitive performance. We showcase the efficacy of our network in tagging top-quark jets in a sea of other light quark and gluon jets.
@@ 1 Introduction
As we enter the high-energy (HE) and high luminosity (HL) phase of the Large Hadron Collider (LHC), the study of jet physics has become more relevant than ever. ‘Fat jet tagging’ is a critical task for identifying the hadronic decays of massive particles, such as the Higgs boson, top quark, or a $W$ or $Z$-boson that are produced with high momentum or ‘boost’ in a particular direction. As a result of this boost, the decay products of these standard model (SM) particles are highly collimated and can be contained within a single large-radius jet, which we call a ‘fat jet’. Scenarios like this can easily occur in various beyond the standard model (BSM) frameworks with vector-like quarks Yang:2025ktj ; Ghosh:2025gdq , leptoquarks Sahoo:2025kdj ; Bhaskar:2024snl ; Ghosh:2025gue , heavy right-handed neutrinos Chakraborty:2018khw ; Padhan:2022fak , supersymmetric particles with $R$-parity violating couplings Baruah:2024wrn ; Sahu:2025uop , to name a few. Among the above-mentioned SM particles, top quark holds a unique status. It happens to be the heaviest fundamental particle we know thus far, and unlike other quarks, it decays before hadronization. The heavy mass of the top quark means its coupling strength with the Higgs boson is the strongest. Hence, precise measurements of top quark properties provide a consistency check of the SM. Even subtle deviations from the SM predictions for top quark interactions could hint at new physics. Moreover, the top quark mass plays a crucial role in determining the stability of the electroweak vacuum. Though the current measurements suggest that our vacuum is ‘metastable’ Hiller:2024zjp ; Degrassi:2012ry ; Buttazzo:2013uya , this conclusion is highly sensitive to the precise values of the top quark measurements Alekhin:2012py ; Elias-Miro:2011sqh ; Espinosa:2015kwx . Identifying top quark jets as fat jets in a sea of other light quark and gluon jets, therefore, is a rather important task.
@@
Machine learning (ML) has revolutionized this field Sahu:2024fzi ; Bhattacherjee:2022gjq ; Mikuni:2025tar by moving beyond traditional methods that relied on high-level, physics-inspired features like jet mass, jet radius, $n$-subjettiness, energy correlation functions, etc Thaler:2010tr ; Thaler:2011gf ; Stewart:2010tn ; Larkoski:2013eya ; Moult:2016cvt ; Komiske:2017aww ; Abdesselam:2010pt ; Kogler:2018hem ; Altheimer:2012mn ; Altheimer:2013yza ; Adams:2015hiv ; Marzani:2019hun ; Larkoski:2017jix . Deep learning models, on the other hand, instead of using simple, hand-crafted features, train themselves to represent the jets as points in a vector space of features best suited for the subsequent classification. The inputs to these models include the four-momentum components of each constituent particle, global features that carry the kinematical properties of the jet, and particle tracking IDs whenever available. From the ML perspective, the jet physics task at hand can be divided into two broad categories: classification and generation. Both of these ‘downstream’ tasks depend on representing the raw, low-level particle or detector data associated with a jet as points in a dense, continuous vector space (feature space), ready to be effectively processed by any ML algorithm. This process, known as ‘embedding/encoding,’ allows the model to learn a more comprehensive and nuanced representation of the substructure of a jet, which can be utilized in subsequent downstream tasks.
@@
A more direct and fundamental approach to embedding is to represent the jet as a point cloud of its constituent particles. Each ‘point’ is a particle with its own set of features, e.g., four momentum components, particle IDs, charges, and tracking information. GNNs Shlomi:2020gdn ; Duarte:2020ngm ; Thais:2022iok are particularly well-suited for this representation. The jet is now treated as a graph where each particle is a ‘node’, and the connections between them are ‘edges’ Komiske:2018cqr ; Qu:2019gqs ; Bogatskiy:2020tje ; Gong:2022lye ; Bogatskiy:2023nnw ; Bogatskiy:2022czk ; Bogatskiy:2023fug . Constituent information can also be represented as a varying length sequence, especially if it is ordered based on a certain property, e.g., by their transverse momentum $p_{T}$. RNNs—originally developed for processing text or speech—can then be used. The advent of the self- and cross-attention mechanisms gave us the famously parallelizable and scalable Transformer architecture DBLP:journals/corr/VaswaniSPUJGKP17 . Originally used to process and generate natural language sequences, these models can be thought of as special variations of GNNs JoshiTrans2GNN , and can process jets either as sequences of constituents or unordered point clouds.
@@
One jarring issue that plagues the existing networks with the highest accuracy in identifying various jets is the computational cost. Achieving high accuracy boils down to extracting maximal information from the jet representation. It can either be gleaned by formidably large networks or can be preprocessed and fed to a smaller network later. The former is the path taken by transformer-based models Qu:2022mxj ; Mikuni:2021pou ; Wu:2024thh and more costly foundational models like Omnilearn(ed) Mikuni:2025tar ; Mikuni:2024qsr ; Bhimji:2025isp . Even these, when striving for higher accuracies, bring separately extracted features to the networks ‘attention’ [^1] Not adding pairwise information to the multi-head attention of the (Par-T) network results in a $1.2\%$ drop in accuracy and a $20-30\%$ drop in background rejection Qu:2019gqs ; Qu:2022mxj .. In the latter approach, when jets are considered as particle clouds or graphs, creating a fully connected graph is prohibitively costly in terms of time and computation, mainly because the pairwise interactions (edge information222i.e., the abstract message-passing, not physical QCD interactions.) are numerous for a jet with a large number of constituents. These challenges are either bypassed by keeping only the nearest neighbor interactions (ParticleNet Qu:2019gqs ), or by representing data in a geometric algebra over space-time, equivariant under physics-motivated (Lorentz) transformations Bogatskiy:2022czk ; Bogatskiy:2023nnw ; Gong:2022lye . The newer architectures combine both of these approaches Spinner:2024hjm ; Brehmer:2024yqw . Even running these state-of-the-art models on single-GPU machines becomes an ordeal, let alone training them from scratch. With the rising awareness of the virtues of Transfer Learning, more models are now pre-trained on larger datasets with multiple classes, which automatically requires significant model complexity. This considerably increases their performance, but the downside is that the cost of retraining increases many times over for specific downstream tasks.
@@
Our objective in this work is to revisit computationally low-cost CNNs, as they implicitly utilize local, pairwise information (albeit spatially, through images) to some extent. Depending on the knowledge that the inclusion of global jet features enhances tagging efficiency Bhattacherjee:2022gjq , we explore the possibility of constructing a lightweight yet efficient network by incorporating global features with the images. Although computationally lightweight, smaller CNNs such as LeNet lenetPaper cannot achieve competitive accuracies. Extracting more information here is equivalent to increasing image resolution and/or model complexity (depth/width). Although CNNs become notoriously computationally costly when scaled up (e.g., ResNet ResNet1 ), recent developments, such as compound scaling EffNetV1 ; EffNetV2 , combined with depthwise and pointwise convolutionsMobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 ; MNASnet , resulted in networks that reach higher accuracies at a fraction of the cost. As all versions of EfficientNet were originally designed for images with higher resolution, we have built lightweight versions of those, while following the same scaling rules. We aim to explore their capabilities with and without the externally provided global features, while using LeNet as a benchmark for comparison throughout.
@@
(a) $32\times 32$
(b) $32\times 32$, Standardized
(c) $32\xrightarrow{Crop}28$, Standardized
@@
(e) $32\times 32$
(f) $32\times 32$, Standardized
(g) $32\xrightarrow{Crop}28$, Standardized
@@
Figure 1: Representative images of two jets for different resolutions and pixel standardization. Although images in our dataset are rendered in RGB, we present them here using the CEI 1931 XYZ color scheme xyzcolorCEI:1931 , for better visual understanding. All images from the top panel are of one representative top jet, and those below are of one QCD jet. The first column shows the jet constituents as a $32\times 32$ image in the $\Delta\eta-\Delta\phi$ plane, as the channels are combined and rendered. The second panel in the same row contains the same image, with the pixels standardized (the mean image of the dataset is subtracted from the original and then divided by the standard deviation image). The third column displays the images in the second, but cropped to $28\times 28$. The fourth column contains the same jets, first rendered in $40\times 40$, then cropped to $32\times 32$.
@@ 2.1 Machine Configuration
All training and measurements, except for the measurement of inference speed, are performed on a single desktop PC with $64GB$ RAM, a $13^{th}$ generation Intel© i9 processor, and a $12GB$ NVIDIA© RTX A2000 graphics processor. The entire analysis is conducted in Mathematica® with the Wolfram Language, which uses the Apache MXNet® framework in the backend. The original EfficientNet architectures (before modification) were taken from the Wolfram Net Repository effnet:Wolfram . To compare the inference speed of CNNs in this work with some external models, inference times were calculated on another machine with an $11GB$ NVIDIA© GeForce GTX 1080-Ti graphics processor.
@@ 2.2 Data: Images from Constituents
We have used a pre-existing dataset [^3] https://syncandshare.desy.de/public.php/dav/files/llbX3zpLhazgPJ6/?accept=zip that contains both top-quark and mixed light-quark-gluon jets. The objective, as mentioned before, is to identify top quark jets as signals and tag them as fat jets amid the sea of other jets. The events were generated with 14 TeV center-of-mass energy using Pythia8 Sjostrand:2014zea followed by detector simulation performed using Delphes deFavereau:2013fsa . The jets were constructed using the anti-kT algorithm Cacciari:2008gp with FastJet Cacciari:2011ma . The pseudorapidity $|\eta_{j}|$ is required to be <2<2 for the fat top quark jet, whereas the radius parameter is chosen to be $0.8$. It is ensured that a matched parton-level top exists within $\Delta R=0.8$ of this signal jet. Moreover, all decay products of the top quark are required to be within $\Delta R=0.8$ of the jet axis. 4-momenta information is stored for the leading $200$ jet constituents. The leading top-quark jets lie within a $p_{T}$ range of 550650550-650 GeV. The total sample consists of one million top jets alongside another million non-top jets, which are treated here as background. The dataset is divided into training / validation / test ($1.2M/0.4M/0.4M$) sets.
@@
We pre-process the data to create 3-channel images with transverse momentum ($p_{T}$), mass ($m$), and energy ($E$) in each channel, respectively. To create each channel from the jet constituents, the hardest constituent is considered to be the center of the channel, and the rest of the image is constructed by binning the remaining constituents in two dimensions: the difference in pseudo-rapidity ($\Delta\eta$) and that in the azimuthal angle ($\Delta\phi$). The paper associated with the original dataset Kasieczka:2019dbj assumes the calorimeter resolution to be $0.04\times 2.25^{\circ}$, where a fat jet with a radius parameter $R=0.8$ can be covered with $40\times 40$ pixels. Fixing the bin sizes thus keeps our jet images at this dimension or less ($32\times 32$). Due to these choices and the nature of the dataset, both $\Delta\eta$ and $\Delta\phi$ axes are retained from 1.2-1.2 to $1.2$, without the loss of any constituents (see Figs (1(a)) and (1(e)) ). Unlike the original paper, we do not fix the axis of the second hardest object; rather, we use an image augmentation layer to randomly flip the image along either axis or both during training. We do not use random cropping, as the jets are constructed objects and can always be centered around the hardest $p_{T}$ constituent for all practical applications. Furthermore, to use N-subjettiness later in our analysis, we also require each jet to have at least 4 constituents, which drops a meager 33 jets from the $201.8K$ dataset.
@@
Standardization of input data has become common in ML, as it has been found to improve the performance of generic neural networks. We perform this standardization by creating a ‘mean’ and a ‘standard deviation’ image using a large number of random samples from the dataset. We respectively subtract from and divide every sample by the values of these arrays. This clearly isolates a fairly circular region in the $\Delta\eta-\Delta\phi$ plane containing the bulk of the jet structure (see Figs (1(b)) and (1(f)) ). During some preliminary runs, it has been consistently observed that using images where the central region covers most of the image field yields better classification performance. Thus, we apply a center-crop to the $32\times 32$ images, making them $28\times 28$ (see Figs (1(c)) and (1(g)) ). A similar size of the central region is obtained for the $40\times 40$ images when they are cropped to $32\times 32$ (see Fig.s (1(d)) and (1(h)) ). We keep more of the field of images than is usual practice in the community, where final images are generally cropped to $R=0.8$ Kasieczka:2017nvn ; Kasieczka:2019dbj ; Macaluso:2018tck . Our crops vary within $0.96\leq R\leq 1.05$.

### v2

@@ Abstract
While Transformer-based and standard Graph Neural Networks (GNNs) have proven to be the best performers in classifying different types of jets, they require substantial computational power. We explore the scope of using a lightweight and scalable version of EfficientNet architecture, along with global features of the jet. The end product is computationally inexpensive but is capable of competitive performance. We showcase the efficacy of our network in tagging top-quark jets in a sea of other light quark and gluon jets. The work also sheds light on the importance of global features for both the accuracy and the apparent redundancy of the network’s complexity.
@@ 1 Introduction
As we enter the high-energy (HE) and high luminosity (HL) phase of the Large Hadron Collider (LHC), the study of jet physics has become more relevant than ever. ‘Fat jet tagging’ is a critical task for identifying the hadronic decays of massive particles, such as the Higgs boson, top quark, or a $W$ or $Z$-boson that are produced with high momentum or ‘boost’ in a particular direction. As a result of this boost, the decay products of these standard model (SM) particles are highly collimated and can be contained within a single large-radius jet, which we call a ‘fat jet’. Scenarios like this can easily occur in various beyond the standard model (BSM) frameworks with vector-like quarks Yang et al. (2025); Ghosh et al. (2025a), leptoquarks Sahoo et al. (2025); Bhaskar and Mitra (2025); Ghosh et al. (2025b), heavy right-handed neutrinos Chakraborty et al. (2019); Padhan et al. (2022), supersymmetric particles with $R$-parity violating couplings Baruah et al. (2025); Sahu et al. (2025), to name a few. Among the above-mentioned SM particles, top quark holds a unique status. It happens to be the heaviest fundamental particle we know thus far, and unlike other quarks, it decays before hadronization. The heavy mass of the top quark means its coupling strength with the Higgs boson is the strongest. Hence, precise measurements of top quark properties provide a consistency check of the SM. Even subtle deviations from the SM predictions for top quark interactions could hint at new physics. Moreover, the top quark mass plays a crucial role in determining the stability of the electroweak vacuum. Though the current measurements suggest that our vacuum is ‘metastable’ Hiller et al. (2024); Degrassi et al. (2012); Buttazzo et al. (2013), this conclusion is highly sensitive to the precise values of the top quark measurements Alekhin et al. (2012); Elias-Miro et al. (2012); Espinosa (2016). Identifying top quark jets as fat jets in a sea of other light quark and gluon jets, therefore, is a rather important task.
@@
Machine learning (ML) has revolutionized this field Sahu et al. (2024); Bhattacherjee et al. (2024); Mikuni and Nachman (2025a) by moving beyond traditional methods that relied on high-level, physics-inspired features like jet mass, jet radius, $n$-subjettiness, energy correlation functions, etc Thaler and Van Tilburg (2011, 2012); Stewart et al. (2010); Larkoski et al. (2013); Moult et al. (2016); Komiske et al. (2018); Abdesselam and others (2011); Kogler and others (2019); Altheimer and others (2012, 2014); Adams and others (2015); Marzani et al. (2019); Larkoski et al. (2020). Deep learning models, on the other hand, instead of using simple, hand-crafted features, train themselves to represent the jets as points in a vector space of features best suited for the subsequent classification. The inputs to these models include the four-momentum components of each constituent particle, global features that carry the kinematical properties of the jet, and particle tracking IDs whenever available. From the ML perspective, the jet physics task at hand can be divided into two broad categories: classification and generation. Both of these ‘downstream’ tasks depend on representing the raw, low-level particle or detector data associated with a jet as points in a dense, continuous vector space (feature space), ready to be effectively processed by any ML algorithm. This process, known as ‘embedding/encoding,’ allows the model to learn a more comprehensive and nuanced representation of the substructure of a jet, which can be utilized in subsequent downstream tasks.
@@
A more direct approach to embedding is to represent the jet as a point cloud of its constituent particles. Each ‘point’ is a particle with its own set of features, e.g., four momentum components, particle IDs, charges, and tracking information. GNNs Shlomi et al. (2020); Duarte and Vlimant (2020); Thais et al. (2022) are particularly well-suited for this representation. The jet is now treated as a graph where each particle is a ‘node’, and the connections between them are ‘edges’ Komiske et al. (2019); Qu and Gouskos (2020); Bogatskiy et al. (2020); Gong et al. (2022); Bogatskiy et al. (2024, 2022, 2023). Constituent information can also be represented as a varying length sequence, especially if it is ordered based on a certain property, e.g., by their transverse momentum $p_{T}$. RNNs—originally developed for processing text or speech—can then be used. The advent of the self- and cross-attention mechanisms gave us the famously parallelizable and scalable Transformer architecture Vaswani et al. (2017). Originally used to process and generate natural language sequences, these models can be thought of as special variations of GNNs Joshi (2025), and can process jets either as sequences of constituents or unordered point clouds.
@@
One jarring issue that plagues the existing networks with the highest accuracy in identifying various jets is the computational cost. Achieving high accuracy boils down to extracting maximal information from the jet representation. It can either be gleaned by formidably large networks or can be preprocessed and fed to a smaller network later. The former is the path taken by transformer-based models Qu et al. (2022); Mikuni and Canelli (2021); Wu et al. (2025) and more costly foundational models like Omnilearn(ed) Mikuni and Nachman (2025a, b); Bhimji et al. (2025). Even these, when striving for higher accuracies, bring separately extracted features to the networks ‘attention’ [^1] Not adding pairwise information to the multi-head attention of the (Par-T) network results in a $1.2\%$ drop in accuracy and a $20-30\%$ drop in background rejection Qu and Gouskos (2020); Qu et al. (2022).. In the latter approach, when jets are considered as particle clouds or graphs, creating a fully connected graph is prohibitively costly in terms of time and computation, mainly because the pairwise interactions (edge information222i.e., the abstract message-passing, not physical QCD interactions.) are numerous for a jet with a large number of constituents. These challenges are either bypassed by keeping only the nearest neighbor interactions (ParticleNet Qu and Gouskos (2020)), or by representing data in a geometric algebra over space-time, equivariant under physics-motivated (Lorentz) transformations Bogatskiy et al. (2022, 2024); Gong et al. (2022). The newer architectures combine both of these approaches Spinner et al. (2024); Brehmer et al. (2025). Even running these state-of-the-art models on single-GPU machines becomes an ordeal, let alone training them from scratch. With the rising awareness of the virtues of Transfer Learning, more models are now pre-trained on larger datasets with multiple classes, which automatically requires significant model complexity. This considerably increases their performance, but the downside is that the cost of retraining increases many times over for specific downstream tasks.
@@
Our objective in this work is to revisit computationally low-cost CNNs, as they implicitly utilize local, pairwise information (albeit spatially, through images) to some extent. Depending on the knowledge that the inclusion of global jet features enhances tagging efficiency Bhattacherjee et al. (2024), we explore the possibility of constructing a lightweight yet efficient network by incorporating global features with the images. Although computationally lightweight, smaller CNNs such as LeNet Lecun et al. (1998) cannot achieve competitive accuracies. Extracting more information here is equivalent to increasing image resolution and/or model complexity (depth/width). Although CNNs become notoriously computationally costly when scaled up (e.g., ResNet He et al. (2016)), recent developments, such as compound scaling Tan and Le (2019, 2021), combined with depthwise and pointwise convolutionsHoward et al. (2017); Sandler et al. (2018); Howard et al. (2019); Qin et al. (2024); Tan et al. (2019), resulted in networks that reach higher accuracies at a fraction of the cost. As all versions of EfficientNet were originally designed for images with higher resolution, we have built lightweight versions of those, while following the same scaling rules. We aim to explore their capabilities with and without the externally provided global features, while using LeNet as a benchmark for comparison throughout.
@@
(a) $35\times 35$
(b) $35\times 35$, Standardized
(c) $35\xrightarrow{Crop}28$, Standardized
@@
(e) $35\times 35$
(f) $35\times 35$, Standardized
(g) $35\xrightarrow{Crop}28$, Standardized
@@
Figure 1: Representative images of two jets for different resolutions and pixel standardization. Although images in our dataset are rendered in RGB, we present them here using the CEI 1931 XYZ color scheme Smith and Guild (1931), for better visual understanding. All images from the top panel are of one representative top jet, and those below are of one QCD jet. The first column shows the jet constituents as a $35\times 35$ image in the $\Delta\eta-\Delta\phi$ plane, as the channels are combined and rendered. The second panel in the same row contains the same image, with the pixels standardized (the mean image of the dataset is subtracted from the original and then divided by the standard deviation image). The third column displays the images in the second, but cropped to $28\times 28$. The fourth column contains the same jets, first rendered in $40\times 40$, then cropped to $32\times 32$.
@@ 2.1 Machine Configuration
All training and measurements, except for the measurement of inference speed, are performed on a single desktop PC with $64GB$ RAM, a $13^{th}$ generation Intel© i9 processor, and a $12GB$ NVIDIA© RTX A2000 graphics processor. The entire analysis is conducted in Mathematica® with the Wolfram Language, which uses the Apache MXNet® framework in the backend. The original EfficientNet architectures (before modification) were taken from the Wolfram Net Repository M. Tan (2020). To compare the inference speed of CNNs in this work with some external models, inference times were calculated on another machine with an $11GB$ NVIDIA© GeForce GTX 1080-Ti graphics processor.
@@ 2.2 Data: Images from Constituents
We have used a pre-existing dataset [^3] https://syncandshare.desy.de/public.php/dav/files/llbX3zpLhazgPJ6/?accept=zip that contains both top-quark and mixed light-quark-gluon jets. The objective, as mentioned before, is to identify top quark jets as signals and tag them as fat jets amid the sea of other jets. The events were generated with 14 TeV center-of-mass energy using Pythia8 Sjöstrand et al. (2015) followed by detector simulation performed using Delphes de Favereau et al. (2014). The jets were constructed using the anti-kT algorithm Cacciari et al. (2008) with FastJet Cacciari et al. (2012). The pseudorapidity $|\eta_{j}|$ is required to be <2<2 for the fat top quark jet, whereas the radius parameter is chosen to be $0.8$. It is ensured that a matched parton-level top exists within $\Delta R=0.8$ of this signal jet. Moreover, all decay products of the top quark are required to be within $\Delta R=0.8$ of the jet axis. 4-momenta information is stored for the leading $200$ jet constituents. The leading top-quark jets lie within a $p_{T}$ range of 550650550-650 GeV. The total sample consists of one million top jets alongside another million non-top jets, which are treated here as background. The dataset is divided into training / validation / test ($1.2M/0.4M/0.4M$) sets.
@@
We pre-process the data to create 3-channel images with transverse momentum ($p_{T}$), mass ($m$), and energy ($E$) in each channel, respectively. To create each channel from the jet constituents, the hardest constituent is considered to be the center of the channel, and the rest of the image is constructed by binning the remaining constituents in two dimensions: the difference in pseudo-rapidity ($\Delta\eta$) and that in the azimuthal angle ($\Delta\phi$). The paper associated with the original dataset Butter and others (2019) assumes the calorimeter resolution to be $0.04\times 2.25^{\circ}$, where a fat jet with a radius parameter $R=0.8$ can be covered with $40\times 40$ pixels. Both $\Delta\eta$ and $\Delta\phi$ axes are retained from 1.2-1.2 to $1.2$ for this dataset, without the loss of any constituents. We keep more of the field of images than is usual practice in the community, where final images are generally cropped to $R=0.8$ Kasieczka et al. (2017); Butter and others (2019); Macaluso and Shih (2018); our images are center-cropped at $R=0.96$. To keep the images as small as possible, we use two sizes in the course of this work: $35\times 35$ (see Figs (1(a)) and (1(e)) ) and $40\times 40$. In the original paper, the axis of the image is rotated in such a way that the second hardest $p_{T}$ object falls on the positive $\Delta\phi$ axis and the third hardest constituent in the first quadrant. We have found out that this results in a lower accuracy for classification; rather, we use an image augmentation layer to randomly flip the image along either axis or both during training. We do not use random cropping, as the jets are constructed objects and can—for all practical applications—be centered around the hardest $p_{T}$ constituent. Furthermore, to use N-subjettiness later in our analysis, we also require each jet to have at least 4 constituents, which drops a meager 33 jets from the $201.8K$ dataset.
@@
Standardization of input data has become common in ML, as it has been found to improve the performance of generic neural networks. We perform this standardization by creating a ‘mean’ and a ‘standard deviation’ image using a large number of random samples from the dataset. We respectively subtract from and divide every sample by the values of these arrays. This clearly isolates a fairly circular region in the $\Delta\eta-\Delta\phi$ plane containing the bulk of the jet structure (see Figs (1(b)) and (1(f)) ). Thus, we apply a center-crop to the $35\times 35$ images, making them $28\times 28$ ($R\sim 0.96$, see Figs (1(c)) and (1(g)) ). A similar size of the central region is obtained for the $40\times 40$ images when they are cropped to $32\times 32$ (see Fig.s (1(d)) and (1(h)) ).